A production operation is making 150 pieces of a product by engaging 5 workers for 300 hours. However, 40 per- cent of the products appear to have various quality problems and the company decides to sell them as seconds at a price of €50 each when the normal product is sold €150 each. What is the produc- tivity of the company? In an effort to improve the situation, a number of initiatives were proposed including a scheme where for every improvement 50 per- cent will be given to workers and the other 50 percent will be held by the company. This resulted in a significant drop in defects, as now only 10 products were faulty out of an output of 130 pieces. What is the new productivity of the company, and what will be the earnings per hour of the workers under the bonus scheme, if the cost per piece is €70 both before and after the scheme?



1
Expert's answer
2021-07-08T13:36:01-0400

The amount of normal units increases from 150×(1 - 0.4) = 90 to 130 - 10 = 120, so all the workers will receive 30×(150 - 70)/2 = 1200, which is 1200/300/5 = £0.8 per unit.